OpenStudy (anonymous):

Will Metal The box plots show student grades on the most recent exam compared to overall grades in the class. two bar graphs shown. The top one is labeled Class. Minimum at 74, Q1 at 78, median at 85, Q3 at 93, maximum at 98. The bottom bar graph is labeled Exam. Minimum at 81, Q1 at 85, median at 93, Q3 at 96, maximum at 99. Which of the following best describes the information about the medians?

OpenStudy (anonymous):

The exam median is only 1-2 points higher than the class median. The exam median is much higher than the class median. The additional scores in the second quartile for the exam data make the median higher. The narrower range for the exam data causes the median to be higher.

OpenStudy (anonymous):

@hoblos

OpenStudy (hoblos):

exam median =93 class median = 85 so what compares those values ?

OpenStudy (anonymous):

so B? the exam median is higher than the class

OpenStudy (hoblos):

yes you are right

OpenStudy (anonymous):

thanks :D could you help me with one more?

OpenStudy (hoblos):

yeah of course i'll try

OpenStudy (anonymous):

The box plots show male and female grades in a U.S. History class. two bar graphs shown. The top one is labeled Male. Minimum at 70, Q1 at 82, median at 88, Q3 at 95. The bottom bar graph is labeled Female. Minimum at 75, Q1 at 78, Q3 at 90, maximum at 92. Which of the following best describes the information about the interquartile ranges? The female interquartile range cannot be known because only three quartiles are showing. The male interquartile range cannot be known because the fourth quartile is missing. The male interquartile range is larger due to the size of the first quartile. The two interquartile ranges are about the same size.

OpenStudy (anonymous):

I think its either a or b im not sure

OpenStudy (hoblos):

interquartile ranges = Q3-Q1 for male : interquartile range = 95-82 =13 for female: interquartile range = 90-78 =12 i think it's d

OpenStudy (anonymous):

ok thanks!

OpenStudy (hoblos):

any time :)